Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by Kavya » Mon Mar 01, 2010 1:09 am
If x = [1,2,3] ; y = [4,5,6] and z = [7,8,9,10].
If numbers are selected at random; whats the probability that xyz will be EVEN?

Answer is 8/9..... how???????
Join the discussion
Source: — Problem Solving |

by bigmonkey31 » Mon Mar 01, 2010 1:36 am
not sure if this is the quickest way.. but you basically need to count.

For each x, you have 1/3 chance of picking each number. For each y, you have a 1/3 chance, and z a 1/4 chance.



Now... we have 3 options with x. And recall that for a product of numbers to be even, you just need ONE of the variables to be even.

These are the combos to get even numbers.

x=1, y=5, z=8,10 P=1/3 x 1/3 x 1/2
x=1, y=4,6, z=7,8,9,10 P=1/3 x 2/3 x 1

x=2, y=4,5,6, z=7,8,9,10 p=1/3 x 1 x 1

x=3, y=5, z =8,10 P=1/3 x 1/3 x 1/2
x=3, y=4,6, z=7,8,9,10 P=1/3 x 2/3 x 1

Add them all up, and you get 16/18 or 8/9
Join the discussion

by sairamGmat » Mon Mar 01, 2010 3:50 am
This one is simple actually...we got to find out combinations of numbers with which product xyz will be odd

x = {1,2,3} y= {4,5,6} z= {7,8,9,10}

here the combinations with which we get odd number are:

x - y - z - product
1 - 5 - 7 - 35
1 - 5 - 9 - 45
3 - 5 - 9 - 135
3 - 5 - 7 - 105

So, there are 4 combinations with which we can get an odd number... As we know the probability of selecting a number from x, y and z are 1/3, 1/3 and 1/4 respectively...

The probability that we can have a odd number are ( 2/3) * (1/3) * (2/4) = 4/36 = 1/9.

So, the probability that we can have an even number is 1-1/9 = 8/9.

Why I wrote 2/3 for x => numbers which can result in ODD product are {1.3}
1/3 for y => numbers which can result in ODD product are {5}
2/4 for z => numbers which can result in ODD product are { 7,9}

That's it... :D
Join the discussion

by November Rain » Mon Mar 01, 2010 4:33 am
Perhaps we coud do a different approach

Instead of computing the probability of an event occuring, perhaps we could do the probability of the opposite event.

For a number to be even in a multiplication, at least one factor must be even, so this means that for a number to be odd no factor can be even.

So, we have the following probabilities:

for X, there are 2/3 of probabilities of comming out odd
for Y, there are 1/3 of probabilities of comming out odd
for Z, there are 1/2 of probabilities of comming out odd

So the answer would be 1 - (2/3 * 1/3 * 1/2) , which is 8/9
Join the discussion

by bigmonkey31 » Mon Mar 01, 2010 11:46 am
November Rain wrote:Perhaps we coud do a different approach

Instead of computing the probability of an event occuring, perhaps we could do the probability of the opposite event.

For a number to be even in a multiplication, at least one factor must be even, so this means that for a number to be odd no factor can be even.

So, we have the following probabilities:

for X, there are 2/3 of probabilities of comming out odd
for Y, there are 1/3 of probabilities of comming out odd
for Z, there are 1/2 of probabilities of comming out odd

So the answer would be 1 - (2/3 * 1/3 * 1/2) , which is 8/9
Definitely the fastest way to do it!
Join the discussion

by analyst218 » Mon Mar 01, 2010 2:15 pm
for sure.
1 - [all X Y Z odd] = even XYZ

1- [2/3 * 1/3 * 2/4] = 1- 1/9 = 8/9

number properties + simple probability
Join the discussion

by Kavya » Thu Mar 04, 2010 7:32 pm
Thanks so so much everyone :-)
Join the discussion

by bpgen » Wed Mar 10, 2010 3:07 am
November Rain wrote:Perhaps we coud do a different approach

Instead of computing the probability of an event occuring, perhaps we could do the probability of the opposite event.

For a number to be even in a multiplication, at least one factor must be even, so this means that for a number to be odd no factor can be even.

So, we have the following probabilities:

for X, there are 2/3 of probabilities of comming out odd
for Y, there are 1/3 of probabilities of comming out odd
for Z, there are 1/2 of probabilities of comming out odd

So the answer would be 1 - (2/3 * 1/3 * 1/2) , which is 8/9

Hey November Rain,
Your posting looks interesting but could you please clarify my doubts below:

I think we could also count probabilities of even nos that coming out for each set as following:

for X, there are 1/3 of probabilities of coming out even
for Y, there are 2/3 of probabilities of coming out even
for Z, there are 2/4=>1/2 of probabilities of coming out even

then XYZ=1/9, do you know anything I'm missing out?
“Ambition is the path to success. Persistence is the vehicle you arrive in.”
Join the discussion

by firdaus117 » Wed Mar 10, 2010 4:26 am
bpgen wrote: Hey November Rain,
Your posting looks interesting but could you please clarify my doubts below:

I think we could also count probabilities of even nos that coming out for each set as following:

for X, there are 1/3 of probabilities of coming out even
for Y, there are 2/3 of probabilities of coming out even
for Z, there are 2/4=>1/2 of probabilities of coming out even

then XYZ=1/9, do you know anything I'm missing out?
You have calculated the probability of xyz being even when all three digits are even.As a result you got lesser value than what it should be actually.Now,you know any single digit if even would turn product into even.That is why it is easier to calculate the opposite-when all three are odd,then only product is odd.
Join the discussion

by GMATSUCKER » Mon Mar 15, 2010 3:30 am
Good one !!! But I'm not sure whether Stuart Kovinsky will like this as this might not be a real GMAT question !!
GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT, GMAT
What's life without GMAT !!!!!!!!
Join the discussion